# Frontier AI’s Impact on the Cybersecurity Landscape

### v3

@@ 4.1 Attacks Targeting Systems
Attack Step 1: Reconnaissance. Real-world attacks increasingly use LLMs and AI agents for active scanning (e.g., satellite capability understanding by Forest Blizzard) and information gathering about victims (e.g., by Salmon Typhoon) [160]. Recent research explores frontier AI in multiple reconnaissance methods: active scanning [221], victim information gathering [56], and open-source database search [189].
@@
Attack Step 2: Weaponization. Malware generation mainly stays at the research level, where real-world incidents mainly report DDoS attacks [20, 230]. Specifically, research shows that LLMs can generate Windows PE malware with high evasion rates against VirusTotal [173] (>>70% for Ransomware, Worm, and DDoS); LLM alignments can be evaded [55, 150]; LLMs are also fine-tuned to generate attacks [201]. For vulnerability identification, LLMs are used in various network and software systems [205, 149, 135, 194, 63, 143]. Real-world attacks also begin employing LLMs for vulnerability detection, including MSDT vulnerability (CVE-2022-30190) [160] and SQLite overflow bugs [93]. Moreover, research shows LLMs can generate PoCs for memory corruption bugs in C/C++ [236, 255, 242]. AI agents can exploit real-world N-day [71, 270] and zero-day [72, 175, 282] vulnerabilities (e.g., SQL/XSS injection, concurrency attacks, remote code execution) and solve CTF challenges [223, 272, 203]. However, the study on exploitation remains mainly at a research level, with only a few public real-world reports [115, 160, 245].
@@
Attack Steps 3-5: Delivery & exploitation & installation. Initial access and installation have been demonstrated in the real world [169, 115], while persistence remains at a research level [199]. Initial access is achieved through compromised accounts or infrastructure [75, 169, 115], with AI assisting in password guessing [107, 238, 104] and bypassing defenses, including intrusion detection, CAPTCHA, and two-factor authentication [181, 57, 160, 53]. LLMs and AI agents can also assist malware configuration and installation [83]. Research works employ frontier AI to maintain persistence across system restarts and credential changes, e.g., AI helps masquerading by disguising malicious activities as benign [199, 23].
@@
Attack Step 6: Commend and control (C&C). Research has demonstrated frontier AI can facilitate privilege escalation [95, 52, 125, 96, 227]. For instance, ChainReactor [52] uses LLM planning to generate exploit chains for privilege escalation. For stable and stealthy C&C channels, attackers can employ LLMs for automated domain generation, creating resilient remote management of compromised systems [199, 11, 23]. Recent research shows LLMs can assist with adaptation [43], allowing self-learning malware to create environment-tailored attacks. Only a few real-world reports exist for credential access in this step [119, 198, 171].
@@
Attack Step 7: Action on objectives. AI-enhanced cyberattacks have proliferated across web, IoT, mobile, and cloud platforms, causing real-world financial losses, privacy leakage, data manipulation, and resource hijacking [20, 109, 230, 115, 75, 169, 48, 269, 78, 30, 260, 211]. A recent attack deploys AsyncRAT malware using LLMs to leak sensitive information [109]. Frontier AI amplifies attacks against web applications and IoT devices in e-commerce, including business logic abuse, API violations, bad bot attacks, and targeted DDoS on retail sites [152, 79]. These attacks disrupt operations, steal data, and exploit digital engagement for financial gain [115]. CrowdStrike reports a 110% rise in cloud-aware cyberattacks and 75% increase in intrusions year-over-year with AI enhancement [48]. Early research explores end-to-end attacks using LLMs or AI agents [250, 239, 229, 211].
@@ 4.2 Attacks Targeting Humans
We find that most human-targeted attacks have demonstrated significant real-world impacts (Figure 2). Studies show a 135% increase in advanced social engineering attacks following ChatGPT’s release from Jan to Feb 2023 [51]. Phishing attacks form a significant portion [32, 70, 25], with voice phishing surging 260% in Q4 2023 compared to Q4 2022 [213]. While studies from 2023 and mid-2024 [32, 68] find AI-generated phishing less effective than human-crafted attacks, recent research [101, 14] demonstrates AI-automated phishing can now match human expert effectiveness.
Malicious AI systems like FraudGPT and WormGPT are actively traded in underground markets [261], with FraudGPT selling for \$200 monthly or \$1,700 annually on dark web platforms and Telegram [197]. These systems analyze social media profiles and public data to craft tailored, convincing phishing messages [27]. AI-powered attacks have significantly increased identity theft and social manipulation [153, 165]. Companies now use AI to generate fraudulent advertising, fake reviews, and celebrity endorsements, prompting the US Federal Trade Commission to implement countermeasures against deceptive practices [210].
Deepfakes represent one of the most serious threats. Hong Kong police documented attackers using AI-driven deepfake technology to impersonate a chief financial officer during a video call, successfully defrauding \$25 million [36]. Child exploitation constitutes another severe AI misuse, including AI-generated child sexual abuse material, automated grooming, and detection evasion techniques [218]. AI-driven bots enable cyberbullying and harassment [106], facilitating emotional manipulation and behavioral control. Some nations reportedly employ AI for cyber espionage and psychological manipulation [192]. Misinformation presents another critical concern, with research demonstrating AI’s growing capabilities to influence human opinions and decisions [182, 18, 13, 82, 47].
@@ 4.3 Proactive and Reactive Defenses
Defense Step 1: Proactive testing include penetration testing and proactive vulnerability detections. Early research works automate penetration testing with LLMs and AI agents [123, 56, 94, 250, 224, 45, 105, 116] and more recent ones focus on optimizing agents’ workflows and tool calls [123, 56]. More and more research leverages AI to facilitate vulnerability detection. This first includes developing customized graph neural networks [281, 252] and uses general LLMs and fine-tuned ones to directly identify vulnerabilities in various software applications [63, 4, 163, 194, 149, 135, 206, 183, 65, 220]. These works report mixed results when comparing AI with traditional static analysis tools (e.g., CodeQL, Snyk, and Fortify) and are all concerned about the errors of AI and limited scalability to repository-level detection involving multiple functions and files. To mitigate these weaknesses and capitalize on LLMs’ strong reasoning capabilities, recent work proposes hybrid workflows that integrate LLMs with static analysis tools for repository-level vulnerability detection [130, 136]. Early evidence shows that AI agents can help with vulnerability detection (e.g., Google Project Zero’s report on a real-world vulnerability: stack buffer underflow in SQLite) [269, 89]. Some research works also improve dynamic fuzzing with frontier AI [112, 205, 254, 168, 156, 58, 273, 275], such as LLM-based pattern/rule/seed generation and RL-based power scheduling [207]. There are also early explorations in using LLMs for hybrid methods that combine static with dynamic program analysis, i.e., HyLLfuzz [155] develops an LLM-based concolic execution that uses LLMs as the solver for finding inputs to specific program branches.
@@
Defense Step 2: Attack detection. Malware detection and network intrusion detection are two applications where ML has been widely used [240, 247, 34, 233, 127, 144]. Recent research shows that frontier AI can address key challenges of traditional AI-based attack detection: reliance on feature engineering and labeled data, and poor generalizability. First, traditional AI models rely heavily on well-designed feature engineering, while deep learning can learn directly from raw inputs. When being used for attack detection, deep learning also does not need feature engineering and can explore large learning spaces beyond human-designed features, potentially leading to more effective solutions.
@@
Second, existing AI-based attack detection requires large amounts of well-labeled data, which is challenging for security data [10, 247]. Transformers can learn general knowledge from vast unsupervised data [231, 59, 29], significantly reducing reliance on large labeled datasets [64, 59]. Recent works show that transformers trained on raw network traffic bytes outperform supervised models trained on handcrafted features in intrusion [91, 140] and malware detection [6, 266].
@@
Third, existing AI-driven attack detection faces generalizability issues with out-of-distribution (OOD) data. Frontier AI models, trained on billions of data points or trillions of tokens, encompass wider distribution ranges during training. This reduces OOD likelihood during inference, which improves the generalizability in intrusion detection [91, 262, 77, 216] and malware detection [6, 266].
@@
Defense Step 3: Triage & forensic. First, for PoC generation, existing works explore improving dynamic fuzzing and symbolic execution with LLMs [241, 99], demonstrating the strong potential of AI agents [242, 166]. For example, recent research [241] proposes using LLMs to translate Python path constraints into Z3 code, which enabled symbolic execution on Leetcode problems with complex control flows and list data structures. FaultiLine [166] introduces an agent that performs dataflow tracing, branch condition reasoning, and test case refinement to generate and validate PoCs for target vulnerabilities. Second, for root cause analysis, traditional methods [90, 162, 50] employ static and dynamic analysis tools, such as reverse execution, backward taint analysis, value-set analysis, and alias analysis. Some recent works explore developing AI agents for root cause analysis [191, 253, 184]. These agents collect additional information (e.g., logs) through tool calling and leverage advanced prompting techniques to improve root analysis performance. Third, for reverse engineering, early works use recurrent networks for function boundary and signature recovery [209, 42]. However, such models have limited generalizability. Recent research [179, 248, 117, 251, 110] explores pre-training or fine-tuning foundation models for binary codes and demonstrates better performance than RNN-based models and traditional rule-based tools in multiple reverse engineering tasks.
@@
Defense Steps 4&5: Remediation development and deployment. Research demonstrates frontier AI’s capability to automatically generate security vulnerability patches [133, 139, 126, 3, 276, 157, 177, 214, 263, 41]. To enhance patching effectiveness, researchers have explored various prompt strategies and built AI agents where LLMs can retrieve contextual information (e.g., caller functions of vulnerable code) [28, 139, 167, 265, 132]. The AIxCC competition demonstrates real-world applications, including patching an Off-by-One bug in SQLite3 [93]. However, no specific work exists on AI-facilitated remediation deployment.
@@ 4.4 Defenses with Provable Guarantees
Formal verification proves whether systems satisfy specific properties. This process is labor-intensive; seL4 required tens of proof-engineer years [102, 103]. Failed verifications require expert refinement, and software updates demand rewriting proofs, significantly increasing human effort. Recent research explores frontier AI to improve formal verification’s automation and scalability. Some works enhance interactive theorem proving [111, 264, 9, 258], e.g., GamePad [111] uses deep learning to synthesize Coq proofs. Others focus on automated program verification, where the typical solution is to model programs in domain-specific formal languages (DSLs), formulate security properties as specifications, and use constraint solvers for verification. Another line of research uses LLMs to generate program invariants, lifting a program to a DSL or supplementary text hints [178, 120, 26, 246, 148, 145]. For example, Loopy [120] leverages LLMs to generate candidate loop invariants and refined invariants via an SMT solver. Lemur [246] uses LLMs to generate high-level program invariants as sub-goals, which are then checked by solvers. LLMLift [26] combines LLMs with verified lifting tools to provide verified code transpilation from the source to DSLs, where LLMs generate lifted codes and proofs for functional equivalence. DafnyBench [148] evaluates SOTA LLMs in generating hints for the Dafny verifier. Other works improve solvers, including using traditional deep learning to enhance SAT and SMT solvers [21, 134] and leveraging LLMs for solver testing and fuzzing [219].

### v4

@@ 4.1 Attacks Targeting Systems
Attack Step 1: Reconnaissance. Real-world attacks increasingly use LLMs and AI agents for active scanning (e.g., satellite capability understanding by Forest Blizzard) and information gathering about victims (e.g., by Salmon Typhoon) [Microsoft]. Recent research explores frontier AI in multiple reconnaissance methods: active scanning [sun2024gptscan], victim information gathering [deng2024pentestgpt], and open-source database search [nvidia_cve_analysis].
@@
Attack Step 2: Weaponization. Malware generation mainly stays at the research level, where real-world incidents mainly report DDoS attacks [ddos, a10ddos2024]. Specifically, research shows that LLMs can generate Windows PE malware with high evasion rates against VirusTotal [pa2023attacker] (>>70% for Ransomware, Worm, and DDoS); LLM alignments can be evaded [deng2023jailbreaker, luo2024jailbreakv]; LLMs are also fine-tuned to generate attacks [setak2024fine]. For vulnerability identification, LLMs are used in various network and software systems [sheng2025large, lu2024grace, li2024llm, sakaoglu2023kartal, ding2024vulnerability, liumake]. Real-world attacks also begin employing LLMs for vulnerability detection, including MSDT vulnerability (CVE-2022-30190) [Microsoft] and SQLite overflow bugs [SQLite3]. Moreover, research shows LLMs can generate PoCs for memory corruption bugs in C/C++ [wan2024cyberseceval, yang2023exploitgen, wang2025cybergym]. AI agents can exploit real-world N-day [fang2024llm, zhang2025bountybench] and zero-day [fang2024teams, patil2024leveraging, zhu2025cvebench] vulnerabilities (e.g., SQL/XSS injection, concurrency attacks, remote code execution) and solve CTF challenges [tann2023using, zhang2024cybench, shao2024empirical]. However, the study on exploitation remains mainly at a research level, with only a few public real-world reports [imperva_news, Microsoft, google_bigsleep].
@@
Attack Steps 3-5: Delivery & exploitation & installation. Initial access and installation have been demonstrated in the real world [openai_report_disrupting, imperva_news], while persistence remains at a research level [deeplocker]. Initial access is achieved through compromised accounts or infrastructure [gmail_account_news, openai_report_disrupting, imperva_news], with AI assisting in password guessing [hitaj2019passgan, wang2023password, securityhero2023] and bypassing defenses, including intrusion detection, CAPTCHA, and two-factor authentication [plesner2024breaking, deng2024oedipusllmenchancedreasoningcaptcha, Microsoft, debicha2023advbot]. LLMs and AI agents can also assist malware configuration and installation [google_adv_misuse]. Research works employ frontier AI to maintain persistence across system restarts and credential changes, e.g., AI helps masquerading by disguising malicious activities as benign [deeplocker, beckerich2023ratgpt].
@@
Attack Step 6: Commend and control (C&C). Research has demonstrated frontier AI can facilitate privilege escalation [happe2024llmshackersautonomouslinux, de2024chainreactor, kujanpaa2021automating, happe2023evaluating, tulla2025alfa]. For instance, ChainReactor [de2024chainreactor] uses LLM planning to generate exploit chains for privilege escalation. For stable and stealthy C&C channels, attackers can employ LLMs for automated domain generation, creating resilient remote management of compromised systems [deeplocker, anderson2016deepdga, beckerich2023ratgpt]. Recent research shows LLMs can assist with adaptation [chung2019availability], allowing self-learning malware to create environment-tailored attacks. Only a few real-world reports exist for credential access in this step [obviam-password-cracking, securityhero-password-cracking, cepa-ai-cyberattacks].
@@
Attack Step 7: Action on objectives. AI-enhanced cyberattacks have proliferated across web, IoT, mobile, and cloud platforms, causing real-world financial losses, privacy leakage, data manipulation, and resource hijacking [ddos, hp_news, a10ddos2024, imperva_news, gmail_account_news, openai_report_disrupting, crowdstrike, zero, gartner_survey, bugcrowd_inside, yang2024seccodeplt, singer2025feasibility]. A recent attack deploys AsyncRAT malware using LLMs to leak sensitive information [hp_news]. Frontier AI amplifies attacks against web applications and IoT devices in e-commerce, including business logic abuse, API violations, bad bot attacks, and targeted DDoS on retail sites [ai-retail, ai-retail2]. These attacks disrupt operations, steal data, and exploit digital engagement for financial gain [imperva_news]. CrowdStrike reports a 110% rise in cloud-aware cyberattacks and 75% increase in intrusions year-over-year with AI enhancement [crowdstrike]. Early research explores end-to-end attacks using LLMs or AI agents [xu2024autoattacker, wang2024sands, usman2024generative, singer2025feasibility].
@@ 4.2 Attacks Targeting Humans
We find that most human-targeted attacks have demonstrated significant real-world impacts (Figure 2). Studies show a 135% increase in advanced social engineering attacks following ChatGPT’s release from Jan to Feb 2023 [darktrace]. Phishing attacks form a significant portion [phishing, eze2024analysis, bethany2024large], with voice phishing surging 260% in Q4 2023 compared to Q4 2022 [voice_phishing]. While studies from 2023 and mid-2024 [phishing, lawenforcggment] find AI-generated phishing less effective than human-crafted attacks, recent research [heiding2024evaluating, spearphishing] demonstrates AI-automated phishing can now match human expert effectiveness.
Malicious AI systems like FraudGPT and WormGPT are actively traded in underground markets [yao2024survey], with FraudGPT selling for \$200 monthly or \$1,700 annually on dark web platforms and Telegram [fraudgpt]. These systems analyze social media profiles and public data to craft tailored, convincing phishing messages [blauth2022artificial]. AI-powered attacks have significantly increased identity theft and social manipulation [theft, openai_report_disrupting_25feb]. Companies now use AI to generate fraudulent advertising, fake reviews, and celebrity endorsements, prompting the US Federal Trade Commission to implement countermeasures against deceptive practices [fakereview].
Deepfakes represent one of the most serious threats. Hong Kong police documented attackers using AI-driven deepfake technology to impersonate a chief financial officer during a video call, successfully defrauding \$25 million [hongkong]. Child exploitation constitutes another severe AI misuse, including AI-generated child sexual abuse material, automated grooming, and detection evasion techniques [sullivan2024ai]. AI-driven bots enable cyberbullying and harassment [cyberbullying], facilitating emotional manipulation and behavioral control. Some nations reportedly employ AI for cyber espionage and psychological manipulation [psy_control]. Misinformation presents another critical concern, with research demonstrating AI’s growing capabilities to influence human opinions and decisions [potter2024hidden, suicide, anthropic_persuasiveness, goldstein2024persuasive, costello2024durably].
@@ 4.3 Proactive and Reactive Defenses
Defense Step 1: Proactive testing include penetration testing and proactive vulnerability detections. Early research works automate penetration testing with LLMs and AI agents [kong2025vulnbot, deng2024pentestgpt, happe2023getting, xu2024autoattacker, temara1maximizing, constantin2023llms, hilario2024generative, isozaki2024towards] and more recent ones focus on optimizing agents’ workflows and tool calls [kong2025vulnbot, deng2024pentestgpt]. More and more research leverages AI to facilitate vulnerability detection. This first includes developing customized graph neural networks [zhou2019devign, xu2017neural] and uses general LLMs and fine-tuned ones to directly identify vulnerabilities in various software applications [ding2024vulnerability, akuthota2023vulnerability, nana2024deep, sakaoglu2023kartal, lu2024grace, li2024llm, shestov2024finetuning, purba2023software, du2024vul, sun2024llm4vuln]. These works report mixed results when comparing AI with traditional static analysis tools (e.g., CodeQL, Snyk, and Fortify) and are all concerned about the errors of AI and limited scalability to repository-level detection involving multiple functions and files. To mitigate these weaknesses and capitalize on LLMs’ strong reasoning capabilities, recent work proposes hybrid workflows that integrate LLMs with static analysis tools for repository-level vulnerability detection [lekssays2025llmxCPG, li2025iris]. Early evidence shows that AI agents can help with vulnerability detection (e.g., Google Project Zero’s report on a real-world vulnerability: stack buffer underflow in SQLite) [zero, repoaudit2025]. Some research works also improve dynamic fuzzing with frontier AI [huang2024large, sheng2025large, yang2023kernelgpt, oliinyk2024fuzzing, meng2024large, deng2023large, zhang2024llamafuzz, zhang2025g2fuzz], such as LLM-based pattern/rule/seed generation and RL-based power scheduling [shi2024bandfuzz]. There are also early explorations in using LLMs for hybrid methods that combine static with dynamic program analysis, i.e., HyLLfuzz [meng2024largehybrid] develops an LLM-based concolic execution that uses LLMs as the solver for finding inputs to specific program branches.
@@
Defense Step 2: Attack detection. Malware detection and network intrusion detection are two applications where ML has been widely used [wang2017adversary, wu2023grim, caviglione2020tight, vinayakumar2019robust, kwon2019survey, liu2025analyzing]. Recent research shows that frontier AI can address key challenges of traditional AI-based attack detection: reliance on feature engineering and labeled data, and poor generalizability. First, traditional AI models rely heavily on well-designed feature engineering, while deep learning can learn directly from raw inputs. When being used for attack detection, deep learning also does not need feature engineering and can explore large learning spaces beyond human-designed features, potentially leading to more effective solutions.
@@
Second, existing AI-based attack detection requires large amounts of well-labeled data, which is challenging for security data [anderson2017machine, wu2023grim]. Transformers can learn general knowledge from vast unsupervised data [vaswani2017attention, devlin2018bert, brown2020language], significantly reducing reliance on large labeled datasets [dosovitskiy2020image, devlin2018bert]. Recent works show that transformers trained on raw network traffic bytes outperform supervised models trained on handcrafted features in intrusion [guthula2023netfound, lin2022bert] and malware detection [al2024exploring, zahan2024shifting].
@@
Third, existing AI-driven attack detection faces generalizability issues with out-of-distribution (OOD) data. Frontier AI models, trained on billions of data points or trillions of tokens, encompass wider distribution ranges during training. This reduces OOD likelihood during inference, which improves the generalizability in intrusion detection [guthula2023netfound, yu2024blockfound, gai2023blockchain, stein2024towards] and malware detection [al2024exploring, zahan2024shifting].
@@
Defense Step 3: Triage & forensic. First, for PoC generation, existing works explore improving dynamic fuzzing and symbolic execution with LLMs [wang2024python, jingxuan2021learch], demonstrating the strong potential of AI agents [wang2025cybergym, nitin2025faultline]. For example, recent research [wang2024python] proposes using LLMs to translate Python path constraints into Z3 code, which enabled symbolic execution on Leetcode problems with complex control flows and list data structures. FaultiLine [nitin2025faultline] introduces an agent that performs dataflow tracing, branch condition reasoning, and test case refinement to generate and validate PoCs for target vulnerabilities. Second, for root cause analysis, traditional methods [guo2019deepvsa, mu2019renn, cui2018rept] employ static and dynamic analysis tools, such as reverse execution, backward taint analysis, value-set analysis, and alias analysis. Some recent works explore developing AI agents for root cause analysis [roy2024exploring, yang2024large, rafi2024enhancing]. These agents collect additional information (e.g., logs) through tool calling and leverage advanced prompting techniques to improve root analysis performance. Third, for reverse engineering, early works use recurrent networks for function boundary and signature recovery [shin2015recognizing, chua2017neural]. However, such models have limited generalizability. Recent research [pei2020xda, xie2024resym, jiang2025beyond, xu2025unleashing, hu2024degpt] explores pre-training or fine-tuning foundation models for binary codes and demonstrates better performance than RNN-based models and traditional rule-based tools in multiple reverse engineering tasks.
@@
Defense Steps 4&5: Remediation development and deployment. Research demonstrates frontier AI’s capability to automatically generate security vulnerability patches [li2025sok, lin2024vulnerabilities, kulsum2024case, ahmed2023better, zhang2024evaluating, meng2024empirical, pearce2023examining, sobania2023analysis, yu2024security, chen2024large]. To enhance patching effectiveness, researchers have explored various prompt strategies and built AI agents where LLMs can retrieve contextual information (e.g., caller functions of vulnerable code) [bouzenia2024repairagent, lin2024vulnerabilities, nong2025appatch, yu2025patchagent, li2025patchpilot]. The AIxCC competition demonstrates real-world applications, including patching an Off-by-One bug in SQLite3 [SQLite3]. However, no specific work exists on AI-facilitated remediation deployment.
@@ 4.4 Defenses with Provable Guarantees
Formal verification proves whether systems satisfy specific properties. This process is labor-intensive; seL4 required tens of proof-engineer years [10years_sel4, labor_sel4]. Failed verifications require expert refinement, and software updates demand rewriting proofs, significantly increasing human effort. Recent research explores frontier AI to improve formal verification’s automation and scalability. Some works enhance interactive theorem proving [huang2018gamepad, yu2023metamath, AlphaProof, yang2024formal], e.g., GamePad [huang2018gamepad] uses deep learning to synthesize Coq proofs. Others focus on automated program verification, where the typical solution is to model programs in domain-specific formal languages (DSLs), formulate security properties as specifications, and use constraint solvers for verification. Another line of research uses LLMs to generate program invariants, lifting a program to a DSL or supplementary text hints [pei2023can, kamath2023finding, bhatia2024verified, wu2023lemur, loughridge2024dafnybench, liu2024propertygpt]. For example, Loopy [kamath2023finding] leverages LLMs to generate candidate loop invariants and refined invariants via an SMT solver. Lemur [wu2023lemur] uses LLMs to generate high-level program invariants as sub-goals, which are then checked by solvers. LLMLift [bhatia2024verified] combines LLMs with verified lifting tools to provide verified code transpilation from the source to DSLs, where LLMs generate lifted codes and proofs for functional equivalence. DafnyBench [loughridge2024dafnybench] evaluates SOTA LLMs in generating hints for the Dafny verifier. Other works improve solvers, including using traditional deep learning to enhance SAT and SMT solvers [balunovic2018learning, li2023g4satbench] and leveraging LLMs for solver testing and fuzzing [sun2023smt].